Translate the following sentence into an equation. Do not solve.
“Triple the difference of a number and fourteen is 60.”


Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi
“Triple the difference of a number and fourteen is 60.”
3(n-14) = 60

3n = 102
n = 34
3(20) = 60

Answer by Alan3354(69443)

n is "the number"
The difference can be n-14 or 14-n
------
3(n-14) = 60
or 3(14-n) = 60